OpenStudy (anonymous):

Write the equation of the line that passes through the points (–1, –4) and (2, 5) in point-slope form.

OpenStudy (ivettef365):

first you need to find the slope using formula

OpenStudy (anonymous):

(2 , 5) (–1, –4) --- 3,9;slope=9/3=3 the equation of the line is: y = mx -mPx + Py , where (Px,Py) is any point that is given and m=slope y = 3x -3(2)+5 y = 3x -6+5 did y = 3x -1 did i do it right ? instruction asks me this --> Part 1: Write the equation of the line that passes through the points (–1, –4) and (2, 5) in point-slope form. Part 2: Using complete sentences, explain whether or not it matters which point is used in the final answer. Also explain why you chose the point you did.

OpenStudy (ivettef365):

yes that is correct y = 3x -1

OpenStudy (anonymous):

Im really stuck on part 2 :((

OpenStudy (ivettef365):

the equation of the line is y = 3x -1, that means that the slope = 3 and the y-int = -1 the line have many points (x,y), so it doesn't really matters what point you used.

OpenStudy (anonymous):

Thank You So Much Best Answer !!

OpenStudy (ivettef365):

you can said that you chose point (2,5) because it was easier for you to deal with positive integers.